# The X-ray properties and structure of A3571 up to R500

### v1

@@ 3.1.3 The radial temperature profile
Thanks to the large FoV of EP-FXT, we are able to explore the outer ICM of A3571 efficiently. Centering on the X-ray peak, we used XSPEC (version 12.14.1) to extract spectra from 11 concentric annular regions (more than 20000 photons were detected in each region) in the 1.0–8.0 keV energy band and conducted a radial analysis. The Ancillary Response Files (ARFs) were generated using the EP-FXT response tool fxtarfgen. The spectra were grouped using grppha with a minimum of 30 counts per bin to ensure the applicability of the statistic. The diffuse ICM thermal emission was modeled using the APEC (version 3.0.9), while Galactic hydrogen absorption was accounted for with the TBABS model. We adopted the solar abundance table from Asplund et al. (2009), with the redshift fixed at 0.039, the abundance fixed at 0.3 and the hydrogen column density fixed at (Willingale et al. 2013). All other parameters were left free during the fit. The radial profiles are shown in Fig. 5. The black points represent the EP-FXT data in this work, respectively. The blue points indicate the XMM-Newton data, while the green points show the ASCA data from Hudaverdi et al. (2005).
@@ 3.2.1 The brightness residual image
To investigate the spatial structure of A3571, we used Sherpa version 4.16.0555https://cxc.cfa.harvard.edu/sherpa/ to subtract an elliptical -model combined with a constant sky background value (see Sect. 3.1.1) from the Fig. 1. The best-fit parameters of the elliptical model are listed in Table 2, and the residual map is shown in Fig. 6.
@@
To enable high-resolution 2D mapping of the ICM thermodynamic properties, we performed an Voronoi tessellation on the X-ray image, growing regions to SNR 60 and refining them with SNR‐weighted Lloyd iterations, yielding 102 adaptive bins. The spectrum extracted from each region was modeled with a TBABS * APEC model in 1.0-8.0 keV, with the same parameter settings described in Sect. 3.1.3. Based on the best-fitting temperature and normalization values, we derived pseudo-pressure and pseudo-entropy maps using and , respectively (Sasaki et al. 2016), is electron density and kT is temperature. Because the normalization obtained from spectral fitting is proportional to the square of the electron density, we express and (Fabian et al. 2006b) (see Fig. 7).
@@
The pseudo-pressure map clearly indicates that the pressure is lower in the northern region and higher in the southern region. In contrast, the pressure distribution derived from XMM-Newton observations appears smooth, showing no prominent localized features, and declines steadily with radius in a similar manner along different azimuthal directions (Hudaverdi et al. 2005). This difference may arise from the different methodologies used to estimate temperature and density. In the case of Hudaverdi et al. (2005), the temperature was estimated from hardness ratios, and the electron density was inferred from the soft-band (0.8–1.6 keV) X-ray surface brightness. While this approach offers advantages in terms of computational efficiency, it may be less capable of resolving subtle structures in dynamically complex regions. In contrast, the EP-FXT analysis employs adaptive binning with a SNR of 60 for spectral fitting, yielding higher spatial resolution maps of temperature and normalization, which better reveal localized pressure features.
@@
At the center, the U-shaped surface brightness deficit exhibits high temperatures and entropy, indicating that it is filled with hot, low-density gas—features reminiscent of X-ray cavities typically driven by AGN feedback. However, A3571 lacks the characteristic radio emission commonly observed in such cavities (Venturi et al. 2002). This absence suggests that AGN feedback is unlikely to be the dominant cause, and that alternative mechanisms—such as mergers or gas sloshing—offer a more plausible explanation for the origin of the U-shaped structure.
@@ 4 DISCUSSION
Although the X-ray morphology of A3571 looks like relaxed, its residual map and 2D thermodynamic maps reveal disturbed structure in the inner region. Considering its galaxy density distribution is also elongated along the north–south direction. To identify possible merging signals, we select two sectors covering the brightness excesses in the northwest and southwest for detailed analysis.
Fig. 6 shows the sectors defined in the northwest and southwest directions, from which the radial surface brightness profiles were extracted and fitted using the best-fit model. In addition, spectra were extracted from each sector and fitted with a TBABS * APEC model, resulting in radial temperature profiles (see Fig. 8).
@@ 5 Conclusion
We performed a systematic X-ray analysis of the galaxy cluster A3571, combining optical observations with X-ray data obtained from the EP-FXT telescope, which features a large FoV and low particle background. We measured the cluster’s outer properties beyond in imaging domain. The main results of this work are summarized as follows:
@@
Our data provides a high-quality image of A3571 within radius . It exhibits a relatively regular morphology. Both the surface brightness and temperature profiles exhibiting typical cool-core characteristics.
@@
Two prominent regions of surface brightness excesses are detected within of the cluster center, located to the north and southwest. The southwestern excess is associated with high-temperature features, whereas the northern excess corresponds to a cooler structure. The northern low-temperature feature is most likely produced by gas sloshing, occurring primarily along the line of sight and triggered by the off-axis passage of a low-mass subcluster.
@@
The temperature distribution of A3571 shows a pronounced north–south asymmetry, consistent with previous XMM-Newton observations. A similar north–south elongation is also evident in the optical galaxy density distribution, suggesting that the cluster’s merger activity likely occurred along this axis.
@@
We detect no thermal emission from the direction of A3572, nor do we find any excess in the density of optical galaxies; therefore, we conclude that this cluster does not exist. In contrast, the galaxy density distribution and X-ray morphology indicate that A3571 is still recovering from a minor merger and is currently in the post-merger phase.
@@ Acknowledgements.
- The Cluster-Merger Shock in 1E 0657-56. arXiv e-prints, pp. arXiv:0803.1803. External Links: Document, 0803.1803, ADS entry Cited by: §1.
@@
- Science objectives of the Einstein Probe mission. arXiv e-prints, pp. arXiv:2501.07362. External Links: Document, 2501.07362, ADS entry Cited by: §1.
@@
- The Instrumental Background of EP/FXT. External Links: 2507.00510, Link Cited by: §2.1.
@@ Appendix A Comparison of temperatures
In addition, the observation of A3571 was carried out during the PV phase in March, whereas the formal calibration campaign was not systematically initiated until April. Consequently, in the absence of finalized calibration products, these data are unavoidably affected by calibration-related systematic uncertainties.

### v2

@@ 3.1.3 The radial temperature profile
Thanks to the large FoV of EP-FXT, we were able to explore the outer ICM of A3571 efficiently. Centering on the X-ray peak, we used XSPEC (version 12.14.1) to extract spectra from 11 concentric annular regions (more than 20000 photons were detected in each region) in the 1.0–8.0 keV energy band and conducted a radial analysis. The ancillary response files were generated using the EP-FXT response tool fxtarfgen. The spectra were grouped using grppha with a minimum of 30 counts per bin to ensure the applicability of the statistic. The diffuse ICM thermal emission was modeled using the APEC (version 3.0.9), while Galactic hydrogen absorption was accounted for with the TBABS model. We adopted the solar abundance table from Asplund et al. (2009), with the redshift fixed at 0.039, the abundance fixed at 0.3 and the hydrogen column density fixed at (Willingale et al. 2013). All other parameters were left free during the fit. The radial profiles are shown in Fig. 5. The black points represent the EP-FXT data in this work, respectively. The blue points indicate the XMM-Newton data, while the green points show the ASCA data from Hudaverdi et al. (2005).
@@ 3.2.1 The brightness residual image
To investigate the spatial structure of A3571, we used Sherpa version 4.16.0999https://cxc.cfa.harvard.edu/sherpa/ to subtract an elliptical -model, centered on the X-ray peak, together with a constant sky background (see Sect. 3.1.1), from Fig. 1. The best-fit parameters of the elliptical model are listed in Table 2, and the residual map is shown in Fig. 6.
@@ 3.2.2 Thermodynamic 2D maps
To enable high-resolution 2D mapping of the ICM thermodynamic properties, we performed an Voronoi tessellation on the X-ray image, growing regions to S/N 60 and refining them with S/N‐weighted Lloyd iterations, yielding 102 adaptive bins. The spectrum extracted from each region was modeled with a TBABS * APEC model in 1.0-8.0 keV, with the same parameter settings described in Sect. 3.1.3. Based on the best-fitting temperature and normalization values, we derived pseudo-pressure and pseudo-entropy maps using and , respectively (Sasaki et al. 2016), is electron density and kT is temperature. Because the normalization obtained from spectral fitting is proportional to the square of the electron density, we express and (see Fig. 7; Fabian et al. 2006b).
@@
The pseudo-pressure map clearly indicates that the pressure is lower in the northern region and higher in the southern region. In contrast, the pressure distribution derived from XMM-Newton observations appears smooth, showing no prominent localized features, and declines steadily with radius in a similar manner along different azimuthal directions (Hudaverdi et al. 2005). This difference may arise from the different methodologies used to estimate temperature and density. In the case of Hudaverdi et al. (2005), the temperature was estimated from hardness ratios, and the electron density was inferred from the soft-band (0.8–1.6 keV) X-ray surface brightness. While this approach offers advantages in terms of computational efficiency, it may be less capable of resolving subtle structures in dynamically complex regions. In contrast, the EP-FXT analysis employs adaptive binning with a S/N of 60 for spectral fitting, yielding higher spatial resolution maps of temperature and normalization, which better reveal localized pressure features.
@@
At the center, the U-shaped surface brightness deficit exhibits high temperatures and entropy, indicating that it is filled with hot, low-density gas—features reminiscent of X-ray cavities typically driven by active galactic nucleus feedback. However, A3571 lacks the characteristic radio emission commonly observed in such cavities (Venturi et al. 2002). This absence suggests that active galactic nucleus feedback is unlikely to be the dominant cause, and that alternative mechanisms—such as mergers or gas sloshing—offer a more plausible explanation for the origin of the U-shaped structure.
@@ 4 Discussion
Although the X-ray morphology of A3571 looks like relaxed, its residual map and 2D thermodynamic maps reveal disturbed structure in the inner region. Considering its galaxy density distribution is also elongated along the north–south direction. To identify possible merging signals, we selected two sectors covering the brightness excesses in the northwest and southwest for detailed analysis.
Figure 6 shows the sectors defined in the northwest and southwest directions, from which the radial surface brightness profiles were extracted and fitted using the best-fit model. In addition, spectra were extracted from each sector and fitted with a TBABS * APEC model, resulting in radial temperature profiles (see Fig. 8).
@@ 5 Conclusion
We performed a systematic X-ray analysis of the galaxy cluster A3571, combining optical observations with X-ray data obtained from the EP-FXT telescope, which features a large FoV and low particle background. We measured the cluster’s outer properties beyond in the imaging domain. The main results of this work are summarized as follows:
@@
Our data provide a high-quality image of A3571 within radius . It exhibits a relatively regular morphology. Both the surface brightness and temperature profiles exhibit typical cool-core characteristics.
@@
Two prominent regions of surface brightness excesses are detected within of the cluster center, located to the north and southwest. The southwestern excess is associated with high-temperature features, whereas the northern excess corresponds to a cooler structure. The northern low-temperature feature is most likely produced by gas sloshing occurring primarily along the line of sight and triggered by the off-axis passage of a low-mass subcluster.
@@
The temperature distribution of A3571 shows a pronounced north–south asymmetry, consistent with previous XMM-Newton observations. A similar north–south elongation is also evident in the optical galaxy density distribution, suggesting that the cluster’s merger activity occurred along this axis.
@@
We detect no thermal emission from the direction of A3572, nor do we find any excess in the density of optical galaxies. This implies that, if A3572 exists, it may host an extremely weak ICM or represent a poor system that is not detectable with the currently available X-ray and optical data. In contrast, the galaxy density distribution and X-ray morphology indicate that A3571 is still recovering from a minor merger and is currently in the post-merger phase.
@@ Acknowledgements.
- The Cluster-Merger Shock in 1E 0657-56: Faster than a Speeding Bullet?. In New Horizons in Astronomy, A. Frebel, J. R. Maund, J. Shen, and M. H. Siegel (Eds.), Astronomical Society of the Pacific Conference Series, Vol. 393, pp. 223. External Links: ADS entry Cited by: §1.
@@
- Science objectives of the Einstein Probe mission. Science China Physics, Mechanics, and Astronomy 68 (3), pp. 239501. External Links: Document, 2501.07362, ADS entry Cited by: §1.
@@
- Characteristics and Modeling of the In-flight Instrumental Background of FXT Onboard Einstein Probe. Research in Astronomy and Astrophysics 25 (11), pp. 115019. External Links: Document, 2507.00510, ADS entry Cited by: §2.1.
@@ Appendix A Comparison of temperatures
In addition, the observation of A3571 was carried out during the performance verification phase in March, whereas the formal calibration campaign was not systematically initiated until April. Consequently, in the absence of finalized calibration products, these data are unavoidably affected by calibration-related systematic uncertainties.